OpenStudy (anonymous):

factor the four term polynomial by grouping x^3+6x^2+8x+48

OpenStudy (akshay_budhkar):

\[x^3+6x^2+8x+48=x^2(x+6)+8(x+6)=(x^2+8)(x+6)\]

OpenStudy (akshay_budhkar):

do you get it?

OpenStudy (anonymous):

no not really can u help me?

OpenStudy (akshay_budhkar):

ok in the first two term i took x^2 common, in the next two terms i took 8 common.. Do you get that part?

OpenStudy (anonymous):

yes

OpenStudy (akshay_budhkar):

ok in the thrid part i take (x+6) common.. so that is how i get that product

OpenStudy (anonymous):

but where do u get (x-6) from?

OpenStudy (akshay_budhkar):

it is x+6

OpenStudy (akshay_budhkar):

i took it common no.. see the first two terms are\[x^3+6x^2\] so taking x^2 common i have\[x^2(x+6)\]

OpenStudy (akshay_budhkar):

you get it?

OpenStudy (anonymous):

yes i do now

OpenStudy (akshay_budhkar):

so is the entire question clear or are u still unclear in some part?

OpenStudy (anonymous):

no i get it now thanks

OpenStudy (akshay_budhkar):

ok happy that you get it =D

OpenStudy (anonymous):

okay wait one more question so how did u end up with (x^2+8)(x+6)